### v2

# RealTraj: Towards Real-World Pedestrian Trajectory Forecasting

Abstract
This paper jointly addresses three key limitations in conventional pedestrian trajectory forecasting: pedestrian perception errors, real-world data collection costs, and person ID annotation costs. We propose a novel framework, RealTraj, that enhances the real-world applicability of trajectory forecasting. Our approach includes two training phases—self-supervised pretraining on synthetic data and weakly-supervised fine-tuning with limited real-world data—to minimize data collection efforts. To improve robustness to real-world errors, we focus on both model design and training objectives. Specifically, we present Det2TrajFormer, a trajectory forecasting model that remains invariant in tracking noise by using past detections as inputs. Additionally, we pretrain the model using multiple pretext tasks, which enhance robustness and improve forecasting performance based solely on detection data. Unlike previous trajectory forecasting methods, our approach fine-tunes the model using only ground-truth detections, significantly reducing the need for costly person ID annotations. In the experiments, we comprehensively verify the effectiveness of the proposed method against the limitations, and the method outperforms state-of-the-art trajectory forecasting methods on multiple datasets. The code will be released at project page.
1 Introduction
The pedestrian trajectory forecasting task aims to predict the future positions of pedestrians based on their past observed states, a critical capability for understanding motion and behavioral patterns [49]. This task plays a crucial role in real-world applications such as social robotics navigation [28, 17], autonomous driving [10], and surveillance systems [14]. Over the years, deep learning models have dominated this domain, with the RNN-based [1, 24, 32], the graph-based [2, 52, 3, 65], and the Transformer-based approaches [22, 21, 53, 50] setting new performance benchmarks on large-scale datasets [76, 45, 33, 48, 4, 71, 40]. However, most existing approaches are evaluated under fixed, well-known benchmarks with ideal experimental conditions, such as perfectly annotated trajectories on the pre-recorded videos. This limits their applicability to real-world scenarios, necessitating further performance enhancements for practical deployment. In this paper, we address three key limitations present in existing approaches, as described below.
Figure 1: Our paper addresses the three limitations in the existing pedestrian trajectory forecasting task. (Top) Pedestrian perception errors can significantly degrade trajectory forecasting performance. (Middle) Real-world data collection necessitates substantial manual effort. (Bottom) Person ID annotations require extensive manual labor.
Pedestrian Perception Errors. Conventional approaches often assume that each pedestrian’s past trajectory is perfectly obtained through upstream perception modules: pedestrian detection and tracking. However, this assumption is unrealistic in real-world scenarios, where obtaining complete trajectories is challenging due to detection errors (e.g., miss-detections and localization errors) and tracking errors (e.g., identity switches). When trajectory forecasting models use imperfect trajectories as input, their performance will degrade significantly [70, 61, 60, 69, 19].
Real-World Data Collection Costs. Conventional trajectory forecasting approaches are typically trained on large-scale datasets and evaluated on domains with data similar to the training conditions (e.g., consistent camera angles and sensor setups). This often overlooks the substantial costs associated with real-world data collection, which often requires significant manual effort. Thus, it is crucial to develop prediction models that can be trained effectively while minimizing real-world data collection efforts.
Person ID Annotation Costs. In addition to the challenge of large-scale real-world data collection, most existing trajectory forecasting models rely on fully supervised training methods. This approach necessitates ground-truth data for future trajectories, including precise pedestrian positions and consistent person identifiers across frames, to minimize the loss between predicted and actual trajectories.
In this paper, we propose a novel trajectory forecasting framework, RealTraj, designed to enhance the real-world applicability of trajectory forecasting by addressing three key limitations. Our approach consists of two training phases—self-supervised pretraining and weakly-supervised fine-tuning—followed by an inference phase. First, we leverage large-scale synthetic trajectory data generated from a simulator [58] to effectively learn trajectory patterns using only synthetic data, reducing the need for extensive real-world data collection (addressing the second limitation). However, relying solely on clean synthetic and real-world data can lead to reduced robustness against real-world errors. To mitigate this, we enhance both the model design and training objectives: we introduce Det2TrajFormer, a trajectory forecasting model designed to be invariant to tracking noise by utilizing past detections as inputs. Additionally, we pretrain the model with multiple self-supervised tasks to improve robustness against perception errors (addressing the first limitation) and enhance forecasting performance solely from detections. Unlike previous trajectory forecasting approaches, we propose training the model exclusively on ground-truth detections during the fine-tuning phase, thereby reducing person ID annotation costs (addressing the third limitation). We further introduce an acceleration regularization term to discourage abrupt changes in acceleration, resulting in smoother and more realistic trajectory forecastings. During inference, Det2TrajFormer takes detection results as input to forecast future trajectories.
In summary, our contributions are threefold: (1) We introduce a novel trajectory forecasting framework with self-supervised pretraining on large-scale synthetic data and weakly-supervised fine-tuning on limited real-world data, enhancing real-world applicability by addressing pedestrian perception errors, data collection costs, and person ID annotation costs. (2) We propose three self-supervised tasks to improve robustness against pedestrian perception errors and enhance trajectory forecasting accuracy using only detections. (3) We reduce person ID annotation costs by training the model solely on ground-truth detections via proposed weakly-supervised loss during the fine-tuning phase.
2 Related Work
2.1 Robust Trajectory Forecasting
Pedestrian trajectory forecasting models aim to predict future positions based on observed trajectories, with deep learning methods demonstrating strong performance due to their representational capabilities [1, 24, 41, 62, 31, 37, 22, 51, 26, 36, 52, 23, 64, 38, 50, 65, 21]. Various architectural designs, such as the RNN-based [1, 24, 32], the graph-based [2, 52, 3, 65], and more recent Transformer-based approaches [22, 21, 53, 50], have been explored to improve forecasting accuracy.
Despite the significant advancements, most trajectory forecasting models operate under idealized conditions, relying on ground-truth past trajectories for training and evaluation while ignoring the impact of imperfect inputs. Recent studies have increasingly recognized the challenges posed by perception module errors [61, 18, 68, 74, 72, 54, 60, 69, 66, 9, 16] and adversarial attacks [73, 5, 75, 29]. This work focuses on enhancing robustness against perception module errors, which are unavoidable in real-world deployments and can significantly degrade prediction accuracy [70, 61, 60, 69, 19]. While some approaches address incomplete observations due to occlusions or limited viewpoints (i.e., detection errors) [18, 68, 72, 54, 66, 9], others tackle tracking errors [70, 61, 60, 74, 16]. However, these methods often utilize architectures and objectives tailored to specific error types, limiting their applicability in diverse real-world scenarios. In this work, we aim to develop a robust model capable of handling both detection and tracking errors.
2.2 Adapting Trajectory Forecasting
Conventional trajectory forecasting models often depend heavily on specific training data domains, overlooking the substantial costs and manual effort required for real-world data collection. To address this challenge, recent research has focused on lightweight methods for adapting pretrained predictors to new data. Some approaches target cross-domain transfer [67, 27, 35], while others emphasize online adaptation [44, 34, 8], continual learning [39, 64, 55], and prompting-based strategies [57]. These methods have demonstrated notable improvements in forecasting performance. However, they often assume access to ground-truth trajectories, which require costly person ID annotations. To reduce this burden, we propose fine-tuning models using only ground-truth detections, thereby eliminating the extensive need for person ID annotations.
2.3 Self-Supervised Learning
Self-supervised learning (SSL) is a promising approach that enables models to learn valuable latent features from unlabeled data. Through pretraining on pretext tasks and pseudo-labels derived from data, followed by fine-tuning on downstream tasks, SSL has facilitated significant advancements in computer vision [25] and natural language processing (NLP) [11]. However, few studies have explored SSL in trajectory forecasting [63, 46, 7, 6, 42, 35, 20]. Applying SSL to trajectory forecasting poses unique challenges, as it typically requires large-scale annotated data for pretraining. Unlike fields such as computer vision and NLP, which benefit from abundant unlabeled data, trajectory forecasting relies on annotated trajectories—often involving costly sensor setups and extensive human annotation—limiting scalability and the potential of SSL. Inspired by advancements in training with synthetic data in image domains [30, 43, 12, 56], we explore self-supervised pretraining using synthetic data to reduce data collection and annotation costs. Additionally, we introduce multiple pretext tasks designed to enhance model robustness against pedestrian perception errors and improve trajectory forecasting performance from detection inputs.
Figure 2: Our proposed framework consists of two training phases and an inference phase. (1) Self-supervised pretraining on synthetic data using multiple pretext tasks. (2) weakly-supervised fine-tuning on real ground-truth detections. (3) Future trajectory inference based solely on detection inputs.
3 RealTraj
3.1 Problem Formulation
We aim to forecast the future trajectory of a target pedestrian in the scene based on detections from observed frames, including the non-targeted pedestrians in the environment, from observed frames. Unlike prior work, our approach skips the tracking step and instead uses detections directly, thereby avoiding the propagation of tracking errors into the forecasting task. Formally, let $Y=(y_{1},y_{2},\dots,y_{T_{pred}})\in\mathbb{R}^{T_{pred}\times 2}$ denote the future trajectory of the target pedestrian over $T_{pred}$ time steps where each position is represented as a tuple $y_{t}=(u_{t},v_{t})\in\mathbb{R}^{2}$, and let $X=(X_{1},X_{2},\dots,X_{T_{obs}})\in\mathbb{R}^{KT_{obs}\times 2}$ represent the set of past detections of $K$ pedestrians over $T_{obs}$ time steps. The detections at time $t$ are defined as $X_{t}=(x_{t}^{1},x_{t}^{2},\dots,x_{t}^{K})\in\mathbb{R}^{K\times 2}$, where each detection is the position of pedestrian $x_{t}^{k}=(u_{t}^{k},v_{t}^{k})\in\mathbb{R}^{2}$ at timestep $t$ [^1] Note that, unlike the conventional trajectory forecasting approaches, our method takes a set of detections as input at each frame. Therefore, it is not guaranteed that $x_{t}^{1}$ and $x_{t+1}^{1}$ correspond to detections of the same pedestrian.. The target pedestrian is one of the $K$ pedestrians in the last observed frame ($t=T_{obs}$). Our goal is to learn a function $\mathcal{F}(X)$ that maps the input detections $X$ to a predicted future trajectory of a target pedestrian $Y$, such that $Y=\mathcal{F}(X)$. In this work, we introduce a novel weakly-supervised trajectory forecasting setting. Large-scale synthetic trajectory data are available during pretraining, and real-world limited detection data are available during fine-tuning.
3.2 Overview
Our framework consists of two training phases—self-supervised pretraining and weakly-supervised fine-tuning—along with an inference phase, as illustrated in Fig. 2. We leverage large-scale synthetic trajectory data from a trajectory simulator [58] to learn effective trajectory patterns, reducing the need for extensive real-world data collection. To ensure real-world robustness, we approach from both the training scheme and the proposed trajectory forecasting model $\mathcal{F}$, Det2TrajFormer. During the pretraining phase, we leverage multiple pretext tasks to improve robustness and forecasting accuracy using large-scale synthetic trajectory data. In the fine-tuning phase, training is conducted solely on ground-truth detections, eliminating the need for costly person ID annotations. During inference, Det2TrajFormer predicts future trajectories based on detection inputs.
Det2TrajFormer.
Det2TrajFormer consists of an embedding layer $\mathcal{G}\in\mathbb{R}^{2\rightarrow d}$, Transformer encoder [59] $\mathcal{E}\in\mathbb{R}^{d\rightarrow d}$, and a forecasting head $\mathcal{P}_{f}\in\mathbb{R}^{d\rightarrow 2}$. Each past detection $x$ is input to the embedding layer to obtain $d$-dimensional embedding along with sinusoidal time index encoding [59]. We concatenate $\mathcal{G}(X)$ with $T_{pred}$ learnable query tokens ($Q\in\mathbb{R}^{T_{pred}\times d}$) and input them into $\mathcal{E}$ to obtain $\hat{H}$ and $\hat{Q}$. $\hat{H}\in\mathbb{R}^{KT_{obs}\times d}$ are output tokens corresponding to the input tokens, $\mathcal{G}(X)$. Finally, the future trajectories $\hat{Y}$ can be predicted from $\hat{Q}$, as specified as follows:
$\begin{split}[\hat{H},\hat{Q}]&=\mathcal{E}(\mathcal{G}(X),Q),\\ \hat{Y}&=\mathcal{P}_{f}(\hat{Q}).\end{split}$ (1)
3.3 Self-Supervised Pretraining Phase
Using synthetically generated trajectories, we train the model on the primary forecasting task while incorporating pretext tasks to enhance performance and improve robustness against detection errors. From the synthetically generated trajectories, we generate a set of past detections $X$ by removing the identity information. Additionally, we simulate the miss-detection and localization errors in the inputs by randomly masking and adding Gaussian noises to $X$, resulting in the corrupted past detections, $X^{C}$. We propose three pretext tasks that aim to reconstruct the removed person ID information from $X$ (person ID reconstruction), reconstruct the masked detections from $X^{C}$ (unmasking), and denoise the $X^{C}$ (denoising). During the pretraining, we additionally employ three heads $\mathcal{P}_{R}\in\mathbb{R}^{d\rightarrow d_{I}}$, $\mathcal{P}_{M}\in\mathbb{R}^{d\rightarrow 2}$, and $\mathcal{P}_{D}\in\mathbb{R}^{d\rightarrow 2}$, where $d$ and $d_{I}$ denote the feature dimension and dimension of person ID embedding.
Trajectory Forecasting Task. The trajectory forecasting task aims to forecast the future trajectory from detections. To improve robustness against detection errors during inference, we forecast the future trajectory $\hat{Y}$ from corrupted detections $X^{C}$ using Eq. 1.
Unmasking Task. The goal of the unmasking pretext task is to encourage the encoder to learn robust features against miss-detections. Specifically, we input $\hat{H}$ obtained from corrupted detection $X^{C}$ by Eq. 1 into the unmasking head $\mathcal{P}_{M}$, as follows:
$\hat{X}_{unmask}=\mathcal{P}_{M}(\hat{H}).$ (2)
Denoising Task. The goal of the denoising pretext task is to encourage the encoder to learn robust features that mitigate localization errors. Similarly, the denoising is formulated as follows:
$\hat{X}_{denoise}=\mathcal{P}_{D}(\hat{H}).$ (3)
Person Identity Reconstruction Task.
The person identity reconstruction pretext task aims to enhance the model’s ability to effectively associate pedestrian instances across observation frames. To achieve this, the network is trained to reconstruct pedestrian identities from detections $X$ by predicting the person identity embeddings. The person identity reconstruction task is formulated as follows:
$\hat{I}=\mathcal{P}_{R}(\hat{H}),$ (4)
where $\hat{I}\in\mathbb{R}^{KT_{obs}\times d_{I}}$ are the reconstructed person ID embeddings for past detections.
Loss.
The model is trained to jointly minimize the losses for the four tasks mentioned above:
$\begin{split}\mathcal{L}=\mathcal{L}_{F}&(\hat{Y},Y)+\alpha\mathcal{L}_{M}(\hat{X}_{unmask},X)\\ &+\beta\mathcal{L}_{D}(\hat{X}_{denoise},X)+\gamma\mathcal{L}_{I}(\hat{I},I),\end{split}$ (5)
where $\mathcal{L}{{}_{F}}$, $\mathcal{L}_{M}$, $\mathcal{L}_{D}$, and $\mathcal{L}_{I}$ represent the trajectory forecasting, unmasking, denoising, and person identity reconstruction losses, respectively. We use the mean square error (MSE) as the loss function between the targets and predictions for each task. The terms $\alpha$, $\beta$, and $\gamma$ are weight hyperparameters.
3.4 Weakly-Supervised Fine-tuning Phase
We fine-tune the model in a weakly-supervised manner by using future ground-truth detections $X_{T_{obs}+1:T_{obs}+T_{pred}}$ instead of trajectories $Y$, thereby reducing ID annotation costs in real-world data. To achieve this, we propose to employ two losses. The first loss is calculated between the predicted future position $\hat{y_{t}}$ and the closest ground-truth detection $d_{t}^{c}$ at each future timestep, summing over the $T_{pred}$ timesteps as follows:
$d_{t}^{c}=\underset{k\in K}{\operatorname{argmin}}\|\hat{y}_{t}-d_{t}^{k}\|_{2},$ (6)
$\mathcal{L}_{W}=\sum_{t=T_{obs}+1}^{T_{obs}+T_{pred}}\|\hat{y}_{t}-d_{t}^{c}\|_{2}.$ (7)
However, this formulation poses a challenge: the predicted position may sometimes deviate from the true trajectory due to its reliance on the closest detection, resulting in unintended oscillations. To address this issue, we introduce an acceleration regularization term, which discourages abrupt changes in acceleration. This promotes smoother transitions and reduces fluctuations in the predicted trajectories. This results in smoother, more realistic trajectory forecasting that aligns with natural motion dynamics. The acceleration regularization term is formulated as follows:
$\mathcal{L}_{Reg}=\sum_{t=T_{obs}+1}^{T_{obs}+T_{pred}-1}\|\hat{y}_{t+1}-2\hat{y}_{t}+\hat{y}_{t-1}\|_{2}.$ (8)
The overall loss function for fine-tuning is defined as:
$\mathcal{L}_{finetune}=\mathcal{L}_{W}+\lambda\mathcal{L}_{Reg},$ (9)
where the terms $\lambda$ is weight hyperparameter.
Refer to caption Refer to caption
Figure 3: Examples of synthetic trajectories generated for self-supervised pretraining.
Table 1: Comparisons with the current state-of-the-art method on JRDB, JTA, ETH-UCY, and SDD under few-shot evaluation. We summarize the ADE metric with different percentages of labeled data. The unit for ADE is meters for JRDB, JTA, and ETH-UCY, and pixels for SDD. Note that the upper bounds for $100\%$ supervised-RealTraj are $0.35$, $0.92$, $0.51$, and $15.99$ for JRDB, JTA, ETH-UCY, and SDD, respectively. The best results are bolded, and the second-best results are underlined.
(a) JRDB
Setting Method $0.1\%$ $0.5\%$ $1\%$ $2\%$ $5\%$
Supervised EqMotion [65]${}_{\text{CVPR'}23}$ 0.45 0.48 0.43 0.60 0.43
Social-Trans [50]${}_{\text{ICLR'}24}$ 0.68 0.48 0.47 0.43 0.43
RealTraj (Ours) 0.43 0.41 0.40 0.38 0.38
Weakly
Supervised
RealTraj (Ours) 0.43 0.41 0.40 0.39 0.38
(b) JTA
Setting Method $0.1\%$ $0.5\%$ $1\%$ $2\%$ $5\%$
Supervised EqMotion [65]${}_{\text{CVPR'}23}$ 1.76 1.68 1.55 1.55 1.55
Social-Trans [50]${}_{\text{ICLR'}24}$ 4.80 3.33 2.91 2.78 2.78
RealTraj (Ours) 1.38 1.25 1.22 1.20 1.07
Weakly
Supervised
RealTraj (Ours) 2.13 1.98 1.68 1.46 1.23
(c) ETH-UCY
Setting Method $0.1\%$ $0.5\%$ $1\%$ $2\%$ $5\%$
Supervised EqMotion [65]${}_{\text{CVPR'}23}$ 0.70 0.56 0.53 0.53 0.53
Social-Trans [50]${}_{\text{ICLR'}24}$ 1.05 0.62 0.55 0.52 0.53
RealTraj (Ours) 0.56 0.55 0.53 0.52 0.52
Weakly
Supervised
RealTraj (Ours) 0.56 0.56 0.56 0.53 0.53
(d) SDD
Setting Method $0.1\%$ $0.5\%$ $1\%$ $2\%$ $5\%$
Supervised EqMotion [65]${}_{\text{CVPR'}23}$ 24.67 19.95 18.53 19.03 17.61
Social-Trans [50]${}_{\text{ICLR'}24}$ 45.02 32.67 23.68 21.80 19.03
RealTraj (Ours) 22.76 20.16 19.22 18.49 17.03
Weakly
Supervised
RealTraj (Ours) 23.49 20.33 19.74 18.99 17.89
4 Experiments
4.1 Datasets
JackRabbot Dataset (JRDB): The JRDB dataset [40], recorded using a social robot, provides diverse real-world pedestrian trajectories in indoor and outdoor settings. We use the train-validation-test split from prior work [50].
Joint Track Auto (JTA) Dataset: The JTA dataset [15] is a large-scale synthetic dataset comprising $256$ training sequences, $128$ validation sequences, and $128$ test sequences set in urban scenarios, generated using a video game engine.
ETH-UCY: The ETH [45] and UCY [33] datasets serve as established benchmarks for human trajectory forecasting, encompassing five scenes. We utilize the leave-one-out validation strategy as described in prior works [24, 37].
Stanford Drone Dataset (SDD): The SDD [48] dataset features over $11,000$ unique pedestrians across $20$ scenes, captured from a bird’s-eye view using a drone. Following the previous trajectory forecasting methods [37], we use the standard setup and train-test split.
TrajImpute: The TrajImpute [9] dataset simulates missing values within ETH-UCY, offering a framework for evaluating trajectory imputation and prediction models.
Following established conventions, we predict $12$ future timesteps based on $9$ past timesteps for the JRDB and JTA datasets, and $12$ future timesteps based on $8$ past timesteps for the ETH-UCY, SDD, and TrajImpute datasets, all at a frame rate of $2.5$ fps.
4.2 Evaluation Metrics
We adopt the widely-used evaluation metrics Average Displacement Error (ADE) and Final Displacement Error (FDE). ADE calculates the average error across all ground truth and predicted positions in the trajectory, while FDE measures the displacement between the endpoints of the ground truth and predicted trajectories.
4.3 Implementation Details
We use the ORCA crowd simulator [58], implemented by [47], to generate synthetic trajectories. Up to $40$ pedestrians are placed in a $15\text{m}\times 15\text{m}$ environment with up to $20$ static primitive obstacles. Examples of generated synthetic trajectories are shown in Fig. 3. We generate $2,000$ sequences for pretraining. We train our model using the Adam optimizer, starting with a learning rate of $1\times 10^{-4}$, which is reduced by a factor of $0.1$ after completing $80\%$ of the total $200$ epochs for pretraining and $20$ epochs for fine-tuning. During the first $100$ epochs of pretraining, the loss weights $\alpha$, $\beta$, and $\gamma$ are set to $1$, 00, and 00, respectively; for the final $100$ epochs, these weights are set to 00, $100$, and $0.1$. A corruption ratio of $30\%$ is applied. Hyperparameters were selected through a coarse-to-fine grid search and step-by-step tuning. The model is trained with a batch size of $16$ on a single NVIDIA TITAN RTX GPU.
Table 2: Comparison with state-of-the-art methods on the JTA, JRDB, ETH-UCY, and SDD. The unit for ADE and FDE is meters for JRDB, JTA, and ETH-UCY, and pixels for SDD. The best results are bolded, and the second-best results are underlined.
Setting Method Input JRDB JTA ETH-UCY SDD
ADE FDE ADE FDE ADE FDE ADE FDE
Supervised Vanilla-LSTM [1]${}_{\text{CVPR'}16}$ Trajectory 0.42 0.83 1.44 3.25 - - - -
Occupancy-LSTM [1]${}_{\text{CVPR'}16}$ 0.43 0.85 1.41 3.15 - - - -
Social-LSTM [1]${}_{\text{CVPR'}16}$ 0.47 0.95 1.21 2.54 0.72 1.54 - -
Social-GAN-ind [24]${}_{\text{CVPR'}18}$ 0.50 0.99 1.66 3.76 0.74 1.54 - -
Transformer [22]${}_{\text{ICPR'}21}$ 0.56 1.10 1.56 3.54 0.54 1.17 - -
Directional-LSTM [32]${}_{\text{T-ITS'}22}$ 0.45 0.87 1.37 3.06 - - - -
Dir-social-LSTM [32]${}_{\text{T-ITS'}22}$ 0.48 0.95 1.23 2.59 - - - -
Trajectron++ [52]${}_{\text{ECCV'}20}$ 0.40 0.78 1.18 2.53 0.53 1.11 - -
Autobots [21]${}_{\text{ICLR'}22}$ 0.39 0.80 1.20 2.70 0.52 1.05 - -
TUTR [53]${}_{\text{ECCV'}20}$ 0.53 1.12 17.38 35.04
EqMotion [65]${}_{\text{CVPR'}23}$ 0.42 0.78 1.13 2.39 0.49 1.03 16.75 33.72
Social-Trans [50]${}_{\text{ICLR'}24}$ 0.37 0.72 0.92 1.84 0.51 1.03 17.96 38.66
RealTraj (Ours) Detection 0.35 0.67 0.92 1.84 0.51 1.03 15.99 31.83
Weakly-Supervised RealTraj (Ours) Detection 0.36 0.71 1.04 2.11 0.52 1.07 16.02 31.98
Refer to caption (b) Miss-Detection Refer to caption (a) Error Localization Refer to caption (c) Identity Switch Refer to caption (d) Combined Error
Figure 4: Comparison of model robustness against various error types on the JRDB dataset, including detection errors (miss-detections and localization errors), tracking errors (identity switches), and their combined impact.
4.4 Few-shot Evaluation
We first demonstrate the effectiveness of RealTraj with limited real-world labeled data. We randomly selected subsets of $0.1\%$, $0.5\%$, $1\%$, $2\%$, and $5\%$ of the dataset, fine-tuned the model on each subset, and evaluated it on the test set on JRDB, JTA, ETH-UCY, and SDD. Since no prior work focuses on a Weakly-Supervised trajectory forecasting method trained with weak detection annotations (denoted as “Weakly-Supervised”), we compare our model with supervised approaches trained with trajectory annotations (denoted as “Supervised”). Specifically, we compare against the current state-of-the-art supervised methods, EqMotion [65] and Social-Transmotion (Social-Trans) [50].
As shown in Tab. 1, our method consistently outperforms these prior methods across all data regimes on JRDB. In the most challenging scenario with only $0.1\%$ of annotations, RealTraj achieves improvements of $4.4\%$, $20\%$, and $4.8\%$ in terms of ADE on JRDB, ETH-UCY, and SDD, respectively, compared to the best-performing previous supervised model, EqMotion. In other data regimes, RealTraj achieves results comparable to supervised methods despite requiring only weak detection annotations during fine-tuning. We also provide results for RealTraj trained in a supervised manner with trajectory annotations (Supervised-RealTraj), which substantially improves over prior supervised methods. Notably, RealTraj closely matches the performance of Supervised-RealTraj on JRDB and SDD across all data regimes. These findings confirm that the proposed model significantly improves performance with limited and weak detection annotations, reducing both data collection and annotation costs, mentioned in Sec. 1.
4.5 Comparison with State-of-the-art
We further compare RealTraj with a wide range of current supervised approaches under a standard $100\%$ labeled evaluation setting, depicted in Tab. 2. The results demonstrate that our framework outperforms all other methods on the JRDB, JTA, and SDD datasets in terms of ADE, surpassing existing approaches by over $4.4\%$ in ADE and $5.2\%$ in FDE on the SDD dataset. Additionally, Supervised-RealTraj achieves the best performance among all methods on JRDB, JTA, ETH-UCY, and SDD in terms of ADE.
4.6 Robustness Evaluation
The robustness of the proposed method against perception errors (detection and tracking errors) is compared with that of the Social-Trans [50] and EqMotion [65], which are the best-performing state-of-the-art method considering both ADE and FDE metrics, as shown in  Tab. 2. We synthetically generated two types of detection errors—miss-detections and localization errors—along with one type of tracking error, specifically identity switches, and combined these errors for evaluation. Miss-detections were simulated by setting past coordinates to zero, while localization errors were introduced by adding Gaussian noise to past coordinates. Identity switches were simulated by swapping identities with nearby pedestrians within a 5-meter radius. Both detection and tracking errors were introduced simultaneously for combined errors. To enable comparison methods to handle missing values in past trajectories, linear interpolation, and extrapolation were applied to fill in missing values before feeding the data into the model.
Fig. 4 (a, b) shows that the performance of comparison methods significantly degrades when detection errors are introduced. In contrast, the proposed method exhibits relatively minor performance degradation, demonstrating robustness against detection errors. Fig. 4 (c) presents a comparison of robustness against tracking errors. While the comparison methods show high sensitivity to tracking errors, our proposed method remains unaffected due to its network architecture, Det2TrajFormer, which forecasts future trajectories directly from detection inputs without relying on tracking information. Lastly, in Fig. 4 (d), we compare robustness against combined detection and tracking errors, where RealTraj consistently outperforms comparison methods by a significant margin across all error ratios.
Table 3: Comparison with the current state-of-the-art method on the Hard-Impute subset of TrajImpute. We report the ADE metric. The unit for ADE is meters. The best results are bolded, and the second-best results are underlined.
Method ETH HOTEL UNV ZARA1 ZARA2 Average
[…]
Supervised-RealTraj 1.11 0.39 0.84 0.59 0.46 0.68
RealTraj 1.17 0.37 0.97 0.78 0.47 0.73
Additionally, we evaluate RealTraj’s robustness against missing data (i.e., miss-detection) using the TrajImpute [9]. Since the combination of the imputation method SAITS [13] and EqMotion achieved the best performance in the TrajImpute benchmark, we adopt SAITS as the imputation method in this evaluation. The results, presented in Tab. 3, show that RealTraj outperforms all comparison methods across all subsets. This includes significant improvements on the challenging HOTEL subset, where the imputation is especially difficult [9]. These findings confirm that the proposed model significantly enhances robustness against perception errors, aligning with our goal of addressing the first robustness limitation described in Sec. 1.
4.7 Ablation Study
Table 4: Ablation study of the proposed pretext tasks on the JRDB dataset. ’F’, ’P’, ’U’, and ’D’ represent forecasting, person ID reconstruction, unmasking, and denoising pretext tasks, respectively. Results are reported for both complete input and corrupted input scenarios with detection errors. The reported metric is ADE.
Main Task Pretext Tasks Complete Detection Errors
F P U D Miss-Detection Error Localization
- - - - 0.78 0.91 0.84
- - - 0.42 0.53 0.70
- 0.39 0.47 0.51
- 0.42 0.52 0.62
- - 0.41 0.51 0.54
- 0.38 0.47 0.51
- 0.38 0.44 0.45
0.36 0.42 0.43
Effectiveness of Pretext Tasks. We first examine the effect of incorporating our proposed pretext tasks during pretraining on the JRDB, as shown in Tab. 4. In addition to using complete inputs, we conducted ablation with inputs containing detection errors to demonstrate the contribution of each pretext task to robustness. Our results show that all pretext tasks enhance trajectory forecasting performance. Specifically, the person identity reconstruction task improves ADE by $7.9\%$ (from $0.39$ to $0.36$). Both the unmasking and denoising tasks also contribute to performance gains, each enhancing ADE by $5.3\%$ (from $0.38$ to $0.36$). Joint training with all three pretext tasks and the main forecasting task improves ADE by $14\%$ (from $0.42$ to $0.36$) compared to pretraining with only the main forecasting task. Overall, the pretraining with our proposed pretext tasks and the main task improves the $54\%$ (from $0.78$ to $0.36$) on ADE compared to the model trained from scratch.
The pretext tasks also contribute to robustness against detection errors. Specifically, the unmasking pretext task improves ADE by $11\%$ when using inputs with miss-detection errors, while the denoising pretext task enhances ADE by $4.4\%$ when using inputs with localization errors. Combining the unmasking and denoising tasks results in further improvements, boosting ADE by $18\%$ (from $0.51$ to $0.42$) for inputs with miss-detections and by $20\%$ (from $0.51$ to $0.43$) for inputs with localization errors. Notably, pretraining is conducted using synthetic trajectories, achieving these results without additional data collection.
Table 5: Ablation study of acceleration regularization weight $\lambda$ on the JTA to examine the impact of the acceleration regularization on overall performance. We selected $\lambda=10$ as the default setting.
Metric $\lambda$
0 1 10 100
ADE 1.16 1.09 1.04 3.67
FDE 2.32 2.18 2.10 5.78
Effectiveness Acceleration Regularization. Hyperparameter acceleration regularization $\lambda$ is used to balance the two terms in Eq. 9. We conducted ablations with different values of acceleration regularization $\lambda$ on JTA in Tab. 5. The results indicate that setting $\lambda=10$ provides the best performance for both ADE and FDE metrics, yielding a $10\%$ improvement in ADE (from $1.16$ to $1.04$) compared to performance without acceleration regularization ($\lambda=0$).
Refer to caption (a) Effect of corruption ratio during training. Refer to caption (b) Effect of the number of synthetic trajectories.
Figure 5: Effect of corruption ratio during training and number of synthetic trajectories.
Effect of Corruption Ratio. We explore the effect of different corruption (adding noise and masking) ratios during pretraining in Fig. 5 (a). Setting the corruption ratios either too low or too high results in suboptimal performance, as the pretext tasks become either too easy or too challenging, limiting their effectiveness for model learning. A moderate corruption ratio yielded the best results.
Effect of the Number of Synthetic Trajectories. Fig. 5 (b) presents the ADE transition when the number of pretraining samples is varied. Our framework shows improvement as the number of data increases.
Refer to caption (a) JRDB Refer to caption (b) JTA
Figure 6: The visualization of predicted trajectories on the JRDB and JTA. The red circle indicates the detection of the target pedestrian in the last observed frame, and the gray circles represent past detections.
5 Qualitative Results
Fig. 6 visualizes the proposed method’s predicted results and the ground-truth future trajectories, highlighting the method’s accuracy in predicting future trajectories in complex scenes only using only past detections.
6 Limitation
Similar to conventional trajectory forecasting approaches, our method can only predict future trajectories for pedestrians detected in the last observed frame. Ideally, predicting future paths for all pedestrians observed in any past frame, even if occluded in the last observed frame, would enhance the applicability of our approach in complex scenes. Addressing this limitation presents an interesting direction for future work.
7 Conclusion
In this paper, we introduced RealTraj, a novel pedestrian trajectory forecasting framework designed to address three key limitations in conventional works. Our approach combines self-supervised pretraining on synthetic data with weakly-supervised fine-tuning on limited real-world data, reducing data collection requirements while enhancing robustness to perception errors. We proposed Det2TrajFormer, a model that leverages past detections to remain robust against tracking noise. By incorporating multiple pretext tasks during pretraining, we further improved the model’s robustness and forecasting accuracy using only detection inputs. Unlike existing methods, our framework fine-tunes only using ground-truth detections, reducing person ID annotation costs. We comprehensively validated its effectiveness against the limitations across multiple datasets, with experimental results demonstrating that our method outperforms state-of-the-art trajectory forecasting approaches.